OpenStudy (anonymous):

Find the area of a circle whose radius is 5 m. A. 78.5 square m B. 314 square m C. 15.7 square m D. 31.4 square m

OpenStudy (anonymous):

just give answer

OpenStudy (texaschic101):

area of a circle = pi times radius squared A = pi * r^2 A = 3.14 * 5^2 A 3.14 * 25 A = 78.5

OpenStudy (anonymous):

thanks